# Risk-Aware Autonomous Driving for Linear Temporal Logic Specifications

### v1

@@ IV Synthesis of human-like risk-aware control
where $r_{th}\in\mathbb{R}$ is a constant threshold.
@@
To sum up, the policy, reward metric, and risk metric are formulated with occupation measures in (11), (12) and (7). Therefore, the control synthesis problem in (8) can be formulated as a linear programming problem using $\beta_{\bar{\pi}}(z,a)\!\geq\!0$ for all $(z,a)\!\in\!Z\!\times\!A$ as decision variables,
@@
(10)$,\;\forall z^{\prime}\in Z.$
This LP problem can be solved using an off-the-shelf solver, Gurobi [40]. If the problem is feasible, an optimal $\bar{\pi}$ can be extracted from the optimized occupation measures using Eq. (11), and then $\bar{\pi}$ can convert to $\pi\!\in\!\Pi$ defined on $\mathcal{M}_{v}$. The obtained policy is trustworthy since its risk metric is in the acceptable range. However, the LP problem may become infeasible sometimes. To address this, a relaxation variable can be introduced into the risk constraint. In practice, the relaxation variable can serve as an indicator of when a human driver should take over control of the vehicle.
V Implementation and Experimental Study
V-A Implementation Framework
The above discussion is based on a MDP model of vehicles. However, to deploy this approach in a real car, we need to determine how to abstract an MDP model from vehicle systems and environments before formulating the proposed method, and how to implement the strategy based on abstract states after solving the LP problem. The complete framework for implementation on autonomous vehicles is illustrated in Fig. 4. More details can be found in the codes of a cute case [41].
@@ V-A Implementation Framework
Vehicle dynamics: In the experiments, we consider the dynamics of vehicles with a discrete-time bicycle model. Specifically, the system state consists of the horizontal position, heading angle, and linear speed of the vehicle, denoted by $[p_{x},p_{y}]$, $\theta$, and $v$, respectively. The control inputs to the vehicle are steering angle and acceleration, denoted by $\phi$ and $a$. Consequently, the system state is $\mathbf{x}(t)=[p_{x}(t),p_{y}(t),\theta(t),v(t)]^{T}\in\mathbb{X}\subset\mathbb{R}^{4}$, and the system input is $\mathbf{u}(t)=[\phi(t),a(t)]^{T}\in\mathbb{U}\subset\mathbb{R}^{2}$, where $\mathbb{X}$ and $\mathbb{U}$ are state space and input space, respectively. The kinematic model of the vehicle is a bicycle model $\mathbf{x}(t+1)\!=\!f(\mathbf{x}(t),\mathbf{u}(t))$ of which the details can be referred to in [42]. To capture the uncertainty of actuation and external factors, a zero-mean Gaussian-distributed disturbance vector, $\mathbf{w}\sim\mathcal{N}(0,\Sigma)$, is considered with the variance $\Sigma\in\mathbb{R}^{4\times 4}$. The vehicle model with uncertainty is denoted by $\mathbf{x}_{k+1}\!=\!f(\mathbf{x}_{k},\mathbf{u}_{k})+\mathbf{w}(t)$.
From continuous dynamics to MDP: To abstract $\mathcal{M}_{v}$ from the vehicle dynamics model with uncertainty, we first discretize the state space $\mathbb{X}$ by gridding, where the finite state set $S$ comprises the centers of the grid cells. For simplicity, this implementation considers only the horizontal positions of vehicles in $\mathcal{M}_{v}$. Similarly, the continuous input space $\mathbb{U}$ is discretized into a finite action set $A$ using appropriate interface functions. In this implementation, the transition probabilities are predefined by experts, which can also be generated autonomously and formally [21]. Additionally, $\mathsf{AP}$ is determined by corresponding propositions in specifications, and the labeling function $\mathcal{L}$ can be extracted from the semantic maps in the perception module.
Control refinement: By solving the LP problem in Eq. (13), an optimal policy, $\pi$ for $\mathcal{M}_{v}$, is obtained such that the next action be determined according to $\pi$. To track the reference state denoted by $\mathbf{x}_{r}\!\in\!\mathbb{X}$ corresponding to the next action, a model predictive controller (MPC) is designed to drive the dynamic model of vehicles, formulated as follows,
@@
where $N$ denotes a horizon number, $X_{0}$ is the current state, $\mathbf{Q}\!\in\!\mathbb{R}^{4\times 4}$ and $\mathbf{R}\!\in\!\mathbb{R}^{2\times 2}$ are diagonal weight matrixes. The first optimal control input $\mathbf{u}_{0}$ is deployed to the vehicle.
V-B Risk Field Visualization
To visualize the risk field intuitively, we first consider a reach-avoid problem over the 2-dimensional positions of the ego vehicle, excluding the product states that include environment states and specification automaton states. Fig. 5 displays the risk field from discounted occupation measures obtained by solving the LP problem in Eq. (13). The red solid line represents the optimal path determined by the optimized risk field. As shown in the figure, we can see that the risk field obtained through the occupation measure aligns with the pattern of the proposed risk field in [24], which verifies the equivalence between the discounted occupation measure and DRF. In addition, the different risk thresholds $r_{th}$ in Eq. (13) will change the shape of the risk field and the corresponding policy, implying different preference for predictive risks.
Figure 5: Risk field extracted from discounted occupation measures. The higher layer is the risk field, while the lower layer is a map of a reach-avoid problem, where the red rectangle is a obstacle and the green rectangle is the target.
V-C Simulation Results
We test the proposed method in three typical scenarios. The video demonstration of three scenarios is accessible at https://youtu.be/0QMtpaREbHE.
@@ V-C Simulation Results
1) Unexpected construction scenario: Real-world traffic frequently encounters abnormal situations that make the original specifications infeasible [11], such as traffic participants violating rules or disruptions of traffic environments. We tested the proposed methods in such a scenario where an unexpected construction zone appears in the forward lane occupied by the ego vehicle, as illustrated in Fig. 6. In this case, the ego vehicle must deviate from a part of the original safety specifications to reach its target. The specification of this scenario for the ego vehicle is ($\square(\neg c\wedge\neg o\wedge\neg s)\wedge\lozenge(t)$, where $c$, $o$, $s$ and $t$ are labels for construction areas, opposite lanes, sidewalks, and target areas. We assign cost values of 10, 5, and 2 to the construction area, sidewalk, and opposite lane, respectively. The generated trajectory is shown in Fig. 6, where each blue rectangle represents the bounding box of the ego vehicle at a previous moment. We can see that the ego vehicle autonomously decides to bypass the construction zone via the opposite lane, which violates the safety specification but with the least violation, and finally reaches the target. This result shows that the proposed risk metric is helpful to yield trajectories that balance different risks and reduce violations as much as possible.
@@
2) Pedestrian crossing scenario: In this scenario, we consider a dynamic environment where a pedestrian is crossing at a crosswalk. The ego vehicle is expected to respond appropriately to the pedestrian’s crossing behavior. To this end, we build an MC model to represent the process of pedestrian crossing behavior. The specification of the ego vehicle is $\square(p\rightarrow\neg c)\wedge\lozenge(t)$, where $c$ and $t$ denote the crosswalk and target zone respectively, and $p$ represents the presence of a pedestrian on the crosswalk. The ego vehicle is expected to autonomously decide whether to stop or proceed based on the real-time situation. To test risk sensitivity, we set different risk thresholds $r_{th}$ in the LP problem in Eq. (13). The simulation results, depicted in Fig. 7, indicate that the ego vehicle will stop at a smaller clearance ahead of the crosswalk as the increase of risk threshold, i.e., $r_{th}=0.1,1,5$. But $r_{th}=10$ is too high to make the ego car account for the collision risk such that the ego car directly hits the pedestrian, which is undesirable. This result demonstrates the parameter $r_{th}$ can be used to tune the conservatism of the synthesized policy.
Figure 7: Pedestrian crossing scenarios: transparent blue vehicles denote stopping points of different risk thresholds.
3) Unprotected turn scenario: We consider a complex intersection scenario that includes a traffic light system and dynamic opponent vehicles, as depicted in Fig. 1. The ego vehicle (blue car) is expected to follow traffic signals and avoid both non-drivable areas and the opponent vehicles (red car). The specification has been introduced in the study case of Sec. II-B. An MC model is designed to represent the switching of traffic lights and the movement of the opponent vehicle. For simplicity, We assume the opponent vehicle always moving forward. The collision risk with the opponent vehicle is calculated by calculating the overlap between risk fields of vehicles as the method used in [28]. The number of discounted occupation measures, i.e., decision variables of the formulated LP problem, is $2880$. With the policy synthesized by the LP problem, the ego vehicle can wait at the boundary of the intersection until the light turns green and adjust its decisions based on the real-time positions of the opponent vehicles. Finally, the ego vehicle reaches the target position and successfully passes the intersection. The whole process is illustrated in Fig. 1 and the attached video, which validates the effectiveness of the proposed method. This case accounts for traffic lights, non-drivable areas, and other dynamic vehicles, which are typical factors in realistic traffic. Therefore, this method has the potential to scale to various scenarios and handle more complex situations.
@@ VI Conclusion
In this paper, we introduce a risk metric for temporal logic specifications using discounting, which can be connected to a typical human-like risk model. The proposed metric accounts for different types of risks in reality. The corresponding control synthesis problem can be solved using an efficient LP problem. Nevertheless, our current approach is still limited by not fully accounting for the interaction between the ego and other vehicles. Also, coarse discretization is introduced for computation efficiency. Future work will focus on addressing these limitations.

### v2

@@ IV Risk-aware control
where $r_{th}\in\mathbb{R}$ is a constant threshold. We first reformulate this optimization problem into a linear programming problem. Afterward, we detail how traffic rules and vehicle dynamics should be integrated and evaluate the risk metric in a traffic scenario.
IV-A Linear Programming Problem
@@
To sum up, the policy, reward metric, and risk metric are formulated with occupation measures in (11), (12) and (7), respectively. Furthermore, the control synthesis problem in (8) can be formulated as a linear programming problem using $\beta_{\bar{\pi}}(z,a)\!\geq\!0$ for all $(z,a)\!\in\!Z\!\times\!A$ as decision variables,
@@
[…]
The optimal policy $\bar{\pi}$ can be extracted from the optimized occupation measures using Eq. (11), and then $\bar{\pi}$ can convert to $\pi\!\in\!\Pi$ defined on $\mathcal{M}_{v}$. This LP problem can be solved using an off-the-shelf solver, Gurobi [35].
To visualize the risk field calculated by the LP problem intuitively, we first consider a reach-avoid problem over the 2-dimensional positions of the ego vehicle, excluding the product states that include environment states and specification automaton states. Fig. 4 displays the risk field with discounted occupation measures obtained by solving the LP problem in Eq. (13). The red solid line represents the optimal path determined by the optimized risk field. As shown in the figure, we can see that the risk field obtained by the occupation measure aligns with the pattern of the proposed risk field in [23], revealing the underlying equivalence between the proposed metric and DRF. In addition, the different risk thresholds $r_{th}$ in Eq. (13) will alter the shape of the risk field and further tune the conservatism of the generated policy.
Figure 4: Risk field extracted from discounted occupation measures. The higher layer is the risk field, while the lower layer is a map of a reach-avoid problem, where the red rectangle is an obstacle and the green rectangle is the target.
IV-B Implementation Framework
The above discussion is based on a MDP model of vehicles. However, to deploy this approach in a real car, we need to abstract an MDP model from vehicle dynamics, as well as determine how to implement high-level policies obtained by solving the LP problem. The complete framework for implementation on autonomous vehicles is illustrated in Fig. 5. The related codes are available in [36].
@@ IV-B Implementation Framework
Vehicle dynamics: In the experiments, we consider the dynamics of vehicles with a discrete-time bicycle model. Specifically, the system state consists of the horizontal position, heading angle, and linear speed of the vehicle, denoted by $[p_{x},p_{y}]$, $\theta$, and $v$, respectively. The control inputs to the vehicle are steering angle and acceleration, denoted by $\phi$ and $a$. Consequently, the system state is $\mathbf{x}(t)\!=\![p_{x}(t),p_{y}(t),\theta(t),v(t)]^{T}\!\in\!\mathbb{X}\!\subset\!\mathbb{R}^{4}$, and the system input is $\mathbf{u}(t)=[\phi(t),a(t)]^{T}\!\in\!\mathbb{U}\!\subset\!\mathbb{R}^{2}$, where $\mathbb{X}$ and $\mathbb{U}$ are state space and input space, respectively. The kinematic model of the vehicle is a bicycle model $\mathbf{x}(t+1)\!=\!f(\mathbf{x}(t),\mathbf{u}(t))$ of which the details can be referred to in [37]. To capture the environmental uncertainty, a zero-mean Gaussian-distributed disturbance vector, $\mathbf{w}\sim\mathcal{N}(0,\Sigma)$, is considered with the variance $\Sigma\in\mathbb{R}^{4\times 4}$. The vehicle model with uncertainty is denoted by $\mathbf{x}(t+1)\!=\!f(\mathbf{x}(t),\mathbf{u}(t))+\mathbf{w}(t)$.
From continuous dynamics to MDP: To abstract $\mathcal{M}_{v}$ from the vehicle dynamics model with uncertainty, we first discretize the state space $\mathbb{X}$ by gridding, where the finite state set $S$ comprises the centers of the grid cells. For simplicity, this implementation considers only the horizontal positions of vehicles in $\mathcal{M}_{v}$. Similarly, the continuous input space $\mathbb{U}$ is discretized into a finite action set $A$ using appropriate interface functions. In this implementation, the transition probabilities are predefined by experts, which can also be generated autonomously and formally [19]. Additionally, the set of atomic propositions $\mathsf{AP}$ is determined by corresponding propositions in specifications, and the labeling function $\mathcal{L}$ can be extracted from the semantic maps in the perception module. Since perception is beyond the scope of this paper, the labels are manually assigned in this implementation.
Control refinement: By solving the LP problem in Eq. (13), an optimal policy, $\pi$ for $\mathcal{M}_{v}$, is obtained such that the next action can be determined according to $\pi$. To track the reference state denoted by $\mathbf{x}_{r}\!\in\!\mathbb{X}$ corresponding to the next action, a model predictive controller (MPC) is designed to drive the dynamic model of vehicles, formulated as follows,
@@
where $N$ denotes a horizon number, $X_{0}$ is the current state, $\mathbf{Q}\!\in\!\mathbb{R}^{4\times 4}$ and $\mathbf{R}\!\in\!\mathbb{R}^{2\times 2}$ are diagonal weight matrixes. The first optimal control input is deployed to the vehicle.
IV-C Pedestrian Crossing Example
We first consider a simple scenario with a dynamic environment where a pedestrian is crossing at a crosswalk. The ego vehicle is expected to respond appropriately to the pedestrian’s crossing behavior. To this end, we build an MC model to represent the uncertain behavior of the pedestrian. The specification of the ego vehicle is $\square(p\!\rightarrow\!\neg\!c)\wedge\lozenge(t)$, where $c$ and $t$ denote the crosswalk and target zone respectively, and $p$ represents the presence of a pedestrian on the crosswalk. The cost of hitting a pedestrian is set to $8$. The ego vehicle is expected to autonomously decide whether to stop or proceed based on the real-time situation. From the risk curve shown in Fig. 6(a), the LP problem in Eq. (13) ensures that the risk always remains below the threshold. To test risk sensitivity, we set different risk thresholds $r_{th}$ for the LP problem. The simulation results, depicted in Fig. 6(b), indicate that the ego vehicle will stop at a smaller clearance ahead of the crosswalk as the increase of risk threshold, i.e., $r_{th}=0.1,1,5$. But $r_{th}=10$ is too high to make the ego car account for the collision risk such that the ego car directly hits the pedestrian, which is undesirable. This result demonstrates the parameter $r_{th}$ can be used to tune the conservatism of the synthesized policy.
(a) Risk curve for $r_{th}\!=\!1$
(b) Stopping points for different risk thresholds.
Figure 6: Pedestrian crossing scenario
In more complex scenarios, the LP problem encounters infeasibility in risky states, requiring a larger $r_{th}$. But a larger $r_{th}$ will lead to unreasonably aggressive behavior in less risky states. As a result, while the formulation in (13) performs well in the pedestrian scenario, it generates unreasonable behavior in complex scenarios, such as unprotected turns in Fig. 1, which cannot be resolved by simply adjusting $r_{th}$. To address this issue, we propose an improved method in the next section.
V Improved risk-aware control and experimental study
Ideally, human drivers can actively minimize real-time risks while avoiding not only catastrophic outcomes but also overly conservative behavior. To achieve such a multi-objective balance, we introduce a soft risk threshold $r_{s}$ and a hard risk threshold $r_{h}$ to manage risk at different levels. This gives the following improved risk-aware control synthesis problem written as an LP optimization,
$\max_{\{\beta_{\bar{\pi}}(z,a)\}}$ $\textstyle\bar{\mathcal{V}}_{\bar{\pi}}(\lozenge G)-K\xi$ (15)
s.t. $\textstyle\sum_{z\in Z}\sum_{a\in A}c(z)\beta_{\bar{\pi}}(z,a)\leq r_{s}+\xi,$
$\xi+r_{s}\leq r_{h},\,\xi\geq 0,$
$\beta_{\bar{\pi}}(z,a)\geq 0,\;\forall(z,a)\in Z\times A,$
[…]
where $\xi$ denotes the relaxation variable for soft risk threshold $r_{s}$ and $K\!>\!0$ is a constant to balance satisfying the co-safety specifications and minimizing the risks of safety specifications. The risk metric is minimized only when the target term forces the risk to exceed the soft threshold. Therefore, $r_{s}$ filters out overly conservative behavior, while $r_{h}$ ensures critical safety.
V-A Experimental Study
We validate the improved risk-aware framework in three different scenarios, using default parameters: risk thresholds $r_{s}\!=\!1$ and $r_{h}\!=\!2$, and discounting factor $\gamma\!=\!0.8$. The related video demonstration is accessible at the link  [^1] https://youtu.be/r5kEMW8oPQE.
1) Pedestrian crossing scenario: The improved method ensures safety and generates behavior with tunable conservatism under the same settings as in Sec. IV-C, demonstrating that the improved LP problem encompasses functionalities of the original one in (13). The maximum and average risk values in the process are listed in Table I.
@@ V-A Experimental Study
2) Unexpected construction scenario: Real-world traffic frequently encounters unexpected situations that make the original specifications infeasible [38], such as traffic participants violating rules or disruptions of traffic environments. We tested the proposed methods in such a scenario where an unexpected construction zone appears in the forward lane of the ego car, as illustrated in Fig. 7. In this case, the ego vehicle must deviate from a part of the original safety specifications to reach its target. The specification of this scenario for the ego vehicle is ($\square(\neg c\wedge\neg o\wedge\neg s)\wedge\lozenge(t)$, where $c$, $o$, $s$ and $t$ are labels for construction areas, opposite lanes, sidewalks, and target areas. We assign cost values of $5$, $3$, and $1$ to the construction area, sidewalk, and opposite lane, respectively. Figure 7 shows the generated trajectory, where the ego vehicle autonomously chooses to bypass the construction zone via the opposite lane with minimal violation and finally reaches the target. This result demonstrates that the improved control method effectively balances different risks and minimizes violations. This advantage arises from incorporating severity into LTL specifications.
@@
3) Unprotected turn scenario: We consider a complex intersection scenario that includes a traffic light system and dynamic opponent vehicles, as depicted in Fig. 1. The ego vehicle (blue car) is expected to follow traffic signals and avoid both non-drivable areas and the opponent vehicles (red car). The specification has been introduced in the study case of Sec. II-B. An MC model is designed to represent the switching of traffic lights and the movement of the opponent vehicle. For simplicity, we assume the opponent vehicle always moving forward. The collision risk with the opponent vehicle is evaluated by calculating the overlap between risk fields of vehicles, following the method described in [24]. With the policy synthesized by the LP problem, the ego vehicle will wait at the boundary of the intersection until the light turns green, adjust its decisions based on the real-time positions of the opponent vehicles, and finally reach the target position. The whole process is illustrated in Fig. 1 and the attached video. This case accounts for traffic lights, non-drivable areas, and other dynamic vehicles, which are typical factors in traffic. Thus, this method can handle complex situations and has the potential to scale to various scenarios.
To examine the influence of different discounting factors, the corresponding risk curves are shown in Fig. 8. While the generated behavior remains identical across different discounting factors, the risk curves in the process are different. Compared to $\gamma\!=\!0.8$, the discounting factor $\gamma\!=\!0.5$ places more emphasis on recent events, causing the risk curves of $\gamma\!=\!0.5$ to gradually increase as the two vehicles approach each other. In contrast, the risk curves for $\gamma\!=\!0.8$ remain flatter because the encounter is perceived earlier. Since placing more focus on recent events lowers perceived risk values, the thresholds are adjusted accordingly.
Figure 8: Risk curves under different discounting factors.
Table I presents the maximum and average risk values with the default discounting and threshold, along with the state number and computation time for the LP problem in (13). The results show that the improved LP problem effectively manages different types of risk and maintains gross risk value within the predefined level throughout the execution process. However, computation time increases significantly as the number of states grows such that a coarse abstraction is adopted. Future work will explore adaptive abstraction and approximation methods to improve computational efficiency.
Scenario 1 Scenario 2 Scenario 3
Maximal risk 0.8 1.46 1.51
Average risk 0.68 0.97 1.01
State number 1440 720 2880
Computation time/ms 20-30 15-20 50-60
TABLE I: Risk value and computation time in the three scenarios.
@@ VI Conclusion
In this paper, we extend a human-like risk metric to account for complex traffic scenarios by integrating LTL specifications. On the other hand, the proposed metric enhances the practicality of temporal logic in traffic tasks by incorporating human-like awareness. Furthermore, a risk-aware control method is improved to balance different types of risks in a human-like manner. Extensive simulation results validate the effectiveness and potential of this framework and illustrate the impact of various parameters. To better handle realistic traffic, future work should incorporate strategic interactions between vehicles, reduce computational costs for finer abstractions, and learn parameters from real-road data instead of manual configuration.